OpenStudy (anonymous):

fan medal and testimonial. help plz what is the instantaneous slope of y= -2/x at x=3? a. -2/9 b. 2/9 c. 2/3 d. -2/3

OpenStudy (anonymous):

a

OpenStudy (astrophysics):

You don't know derivatives yet right

OpenStudy (anonymous):

right.

OpenStudy (astrophysics):

Ok so remember that definition I showed you a while back? \[\lim_{h \rightarrow 0} \frac{ f(a+h)-f(a) }{ h }\]

OpenStudy (astrophysics):

We are just using \[\lim_{h \rightarrow 0} \frac{ f(x+h)-f(x) }{ h }\]

OpenStudy (anonymous):

what is h? I think its 2.

OpenStudy (astrophysics):

That's the definition of the instantaneous rate of change, you have to plug in f(3), how are you getting 2? Guessing isn't going to help on tests.

OpenStudy (anonymous):

I got the 2 from the equation itself y= -2/x. im asking because I am looking for h I know I have to plug in 3 but h has me confused.

OpenStudy (astrophysics):

It's the definition, you must've learned about the slope of a tangent line, and drawing it out the graphs we basically represent \[\Delta x = h\] from \[\frac{ \Delta y }{ \Delta x }\] that's the simplest way I can put it

OpenStudy (anonymous):

ok my answer is 2/3